Determine the number of ordered pairs of positive integers (a,b) such that a and b are both divisors of 1260 and a/b
Answer by greenestamps(13203)



The prime factorization of 1260 is 1260 = 2*2*3*3*5*7

One of the factors of 1260 is 20 = 2*2*5; 1260/20 = 63.

If a=20, then b can be any number of the form 20n, where n is any factor of 63.

63 = 3*3*7; the number of factors of 63 is (2+1)(1+1) = 3*2 = 6.

So there are 6 ordered pairs of positive integers (20,b) for which 20 and b are both divisors of 1260 and a divides b: (20,20), (20,60), (20,140), (20,180), (20,420), and (20,1260).

Perform a similar analysis for each value of a:"

    +-------------------------------------------+
    |   |          |             |  number of   |
    |   |          |             |  possible    |
    |   |          | factors     |   values     |
    | a | n=1260/a |   of n      |   for b      |
    +-------------------------------------------+

      1     1260     2*2*3*3*5*7   3*3*2*2 = 36
      2      630     2*3*3*5*7     2*3*2*2 = 24
      3      420     2*2*3*5*7     3*2*2*2 = 24
      4      315     3*3*5*7       3*2*2   = 12
      5      252     2*2*3*3*7     3*3*2   = 18
      6      210     2*3*5*7       2*2*2*2 = 16
      7      180     2*2*3*3*5     3*3*2   = 18
      9      140     2*2*5*7       3*2*2   = 18
     10      126     2*3*3*7       2*3*2   = 12
     12      105     3*5*7         2*2*2   =  8
     14       90     2*3*3*5       2*3*2   = 12
     15       84     2*2*3*7       3*2*2   = 12
     18       70     2*5*7         2*2*2   =  8
     20       63     3*3*7         3*2     =  6
     21       60     2*2*3*5       3*2*2   = 12
     28       45     3*3*5         3*2     =  6
     30       42     2*3*7         2*2*2   =  8
     35       36     2*2*3*3       3*3     =  9
     36       35     5*7           2*2     =  4
     42       30     2*3*5         2*2*2   =  8
     45       28     2*2*7         3*2     =  6
     60       21     3*7           2*2     =  4
     63       20     2*2*5         3*2     =  6
     70       18     2*3*3         2*3     =  6
     84       15     3*5           2*2     =  4
     90       14     2*7           2*2     =  4
    105       12     2*2*3         3*2     =  6
    126       10     2*5           2*2     =  4
    140        9     3*3           3       =  3
    180        7     7             2       =  2
    210        6     2*3           2*2     =  4
    252        5     5             2       =  2
    315        4     2*2           3       =  3
    420        3     3             2       =  2
    630        2     2             2       =  2
   1260        1                           =  1
                             ------------------
                                total:      330


ANSWER: There are 330 ordered pairs (a,b) in which a and b are both divisors of 1260 and a divides b.